Thanks in advance!

Suppose a simple random sample of size n 75 is obtained from a population whose size is N-20,000 and whose population proportion with a specified characteristic is p 0.6. Complete parts (a) through (c) below. (a) Describe the sampling distribution of p. Choose the phrase that best describes the shape of the sampling distribution below. ○ A. Not normal because n 0.05N and np(1-p) OB. Approximately normal because n s0.05N and np(1-p)<10 yC. Approximately normal because n 0.05N and np(1 p)2 10 0 D. Not normal because n 0.05N and np(1-p) < 10 2 10 Determine the mean of the sampling distribution of p Round s one declps
Answer #1

Answer

we know that the formula for the mean of sampling distribution is given as

n * D

it is given that the sample size is given as n = 75 and population proportion is p = 0.6

setting the given values, we get

n * D = 75*0.6 = 45.0

So, the required sampling mean is 45.0